A wire has a length of 2.0 m and a resistance of 5.0 Ω. Find the electric field existing inside the wire if it carries a current of 10 A.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. l = 2 m, R = 5 Ω, I = 10 A, E = ? V = iR = 10 * 5 = 50 V E = V/l = 50/2 = 25 V/m.